A bag contains 10 red balls, 4 blue balls, and 6 yellow balls. If a ball is picked at random from the bag, what is the probability that the ball is yellow?
1
Expert's answer
2021-08-16T11:20:27-0400

Total number of balls = 10+4+6=20

P(yellow) "= \frac{6}{20}= 0.3"